OpenStudy (anonymous):

How do I solve 4 - (-2) =

OpenStudy (anonymous):

6

OpenStudy (anonymous):

when you are subtracting a negative, you need to alter your equation to solve it. subtracting a negative is equivalent to adding a positive. so in this situation its changed to 4 + 2 =. do you understand?

OpenStudy (texaschic101):

a negative times a negative is a positive. 4 - (-2) = 4 + 2 = 6